find the probability of rolling a fair die and getting a number >4.
Answer by stanbon(57361)

find the probability of rolling a fair die and getting a number >4.
---
That means getting 5 or 6.
P(5 or 6) = P(5)+p(6) = 2*1/6) = 1/3
==========================================
Cheers,
Stan H.
==========